Two POSITIVE NUMBERS HAVE the ratio 2:3. THE LARGER NUMBER IS 30 MORE THAN ONE-HALF OF THE SMALLER NUMBER. FIND THE NUMBER.
Answer by venugopalramana(3286)

Two POSITIVE NUMBERS HAVE the ratio 2:3. THE LARGER NUMBER IS 30 MORE THAN ONE-HALF OF THE SMALLER NUMBER. FIND THE NUMBER.
LET THE 2 NUMBERS BE 2X,3X
ONE HALF OF SMALLER NUMBER = 2X/2=X
30 MORE ...=X+30
3X=X+30
2X=30
X=15
THE 2 NUMBERS ARE
15*2=30 AND 15*3=45